A positive charge, q = +35 nC, is on the y-axis at y = +5.00 cm. (a) Determine the magnitude

and direction of the electric field at the origin. (b) What will be the magnitude and direction

of the electric field at the origin if the charge is -20 nC?


1
Expert's answer
2022-05-16T13:41:00-0400

(a)

"E=k\frac{q}{r^2}"

"E=9*10^9*\frac{35*10^{-9}}{0.05^2}=1.26*10^{5}\:\rm N/C"

Field is directed toward negative x-axis.

(b)

"E=9*10^9*\frac{20*10^{-9}}{0.05^2}=7.2*10^{4}\:\rm N/C"

Field is directed toward positive x-axis.